What age is a boy now if his mother is 12 times his age and in 20 years she will be twice his age?
Answer by alka001(14)

let ptesent age of boy be x yrs
then present age of mother is 12x yrs
after 20 yrs
age of boy will be x+20 yrs
so age of mother will be
12x+20=2(x+20)
solving
12x+20=2x=40
12x-2x=40-20
10x=20
x=2
so prest age boy=2 yrs
present age mother =12*2=24 yrs